# Strichartz estimates for the generalized Zakharov-Kuznetsov equation on ×$\mathbb{R}\times\mathbb{T}$ and applications

### v1

@@ Abstract.
In this article, we address the Cauchy problem associated with the $k$-generalized Zakharov-Kuznetsov equation posed on ×$\mathbb{R}\times\mathbb{T}$. By establishing an almost optimal linear $L^{4}$-estimate, along with a family of bilinear refinements, we significantly lower the well-posedness threshold for all $k\geq 2$. In particular, we show that the modified Zakharov-Kuznetsov equation is locally well-posed in $H^{s}(\mathbb{R}\times\mathbb{T})$ for all $s>\frac{3}{8}$.
@@ 1. Introduction
$E_{\pm}(u(t))\coloneqq\frac{1}{2}\int_{\mathbb{R}\times\mathbb{T}}\lvert\nabla u(t,x,y)\rvert_{2}^{2}\mp\frac{2}{k+2}(u(t,x,y))^{k+2}\mathrm{d}(x,y)$
@@
In the case $k=1$, the foundational work was carried out by Linares, Pastor, and Saut [23], who established local well-posedness in $H^{s}(\mathbb{R}\times\mathbb{T})$ for all $s\geq\frac{3}{2}$. By developing a bilinear estimate that allows for a gain of up to $\frac{1}{2}-$ derivatives, Molinet and Pilod [27] were subsequently able to prove global well-posedness for all $s\geq 1$. Based on a refinement of said estimate, Osawa [25] succeeded in lowering the threshold for local well-posedness to $s>\frac{9}{10}$, and, in joint work with Takaoka [26], further established global existence for all $s>\frac{29}{31}$ by employing the $I$-method. More recently, Cao-Labora [4] developed a further refinement of said bilinear estimate, carefully adapted to the resonance function. By invoking geometric arguments, he was then able to improve the local well-posedness result to $s>\frac{3}{4}$ (or $s>\frac{1}{2}$ under an additional low-frequency condition), and in both cases established optimality up to the endpoint. For the case $k\geq 2$, Farah and Molinet [7] proved the linear $L^{4}$-estimate
@@
The starting point of this article will be the proof of several new linear and bilinear estimates of Strichartz type. In particular, we wish to highlight the following linear $L^{4}$-estimate, which we will furthermore show to be optimal up to the loss of an $\varepsilon$-derivative.
@@
and let $s(k)>s_{0}(k)$ be arbitrary. Then for every $u_{0}\in H^{s(k)}(\mathbb{R}\times\mathbb{T})$, there exists a $\delta=\delta(\lVert u_{0}\rVert_{H^{s_{0}(k)+}(\mathbb{R}\times\mathbb{T})})>0$ and a unique solution
$u\in X_{s(k),\frac{1}{2}+}^{\delta}(\varphi)$
to (1). Moreover, for every $\tilde{\delta}\in(0,\delta)$, there exists a neighborhood $\mathcal{U}\subseteq H^{s(k)}(\mathbb{R}\times\mathbb{T})$ of $u_{0}$ such that the flow map
$S:H^{s(k)}(\mathbb{R}\times\mathbb{T})\supseteq\mathcal{U}\rightarrow X_{s(k),\frac{1}{2}+}^{\tilde{\delta}}(\varphi),\ u_{0}\mapsto S(u_{0})\coloneqq u$
@@
Subsequently, in the case $k=2$ (which corresponds to the modified Zakharov-Kuznetsov equation), we will carry out a detailed case-by-case analysis using the bilinear refinements of (2) and the resonance function, and push the local well-posedness threshold in this case down to $s>\frac{3}{8}$.
@@
Let $s>\frac{3}{8}$ be arbitrary. Then for every $u_{0}\in H^{s}(\mathbb{R}\times\mathbb{T})$, there exists a $\delta=\delta(\lVert u_{0}\rVert_{H^{\frac{3}{8}+}(\mathbb{R}\times\mathbb{T})})>0$ and a unique solution
@@ 2.1. Notation
If $A$ is a measurable subset of $\mathbb{R}\times\mathbb{Z}$, $\mathbb{R}\times\mathbb{R}\times\mathbb{Z}$ or $\mathbb{R}^{n}$, then $\lvert A\rvert$ will denote its Haar measure. For an arbitrary number $x\in\mathbb{R}$, $x+$ wil denote a number slighty greater than $x$, while $x-$ [^1] In the context of interpolation, the notation $\infty$-∞ - will be used to indicate a number that is very large, but still finite. will denote a number slightly smaller than $x$. If we have two real numbers, $x$ and $y$, then their maximum will be denoted by $x\lor y$ and their minimum by $x\land y$. If $x$ and $y$ are both positive, then $x\lesssim y$ will mean that there exists some constant $c>0$, such that $x\leq cy$ holds. In the case that the constant $c$ can be chosen particularly close to 00, we will write $x\ll y$ and if both $x\lesssim y$ and $y\lesssim x$ hold, we will write $x\sim y$. Furthermore, we will denote the absolute value of a real number $x$ by $\lvert x\rvert$, while for a vector $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, the Euclidean norm will be denoted by $\lvert x\rvert_{2}$. Using the Euclidean norm, we define the Japanese brackets as $\langle x\rangle\coloneqq(1+\lvert x\rvert_{2}^{2})^{\frac{1}{2}}$. For an admissible function $f=f(t,x,y)$, where $(t,x,y)\in\mathbb{R}^{2}\times\mathbb{T}$, its space-time Fourier transform will be denoted by $\hat{f}$, and its inverse space-time Fourier transform by $\check{f}$. If we are interested in the partial Fourier transform of such a function, we indicate this by using subscripts, e.g., in $\mathcal{F}_{x}f$ and $\mathcal{F}_{xy}^{-1}f$. We use the Fourier transform to define the Bessel and Riesz potential operators of order $-s$, where $s$ is a real number:
@@
$R(\xi,q,\xi_{1},q_{1})\coloneqq\varphi(\xi,q)-\varphi(\xi_{1},q_{1})-\varphi(\xi-\xi_{1},q-q_{1}).$
@@ 2.2. Function spaces
where $\varphi(\xi,q)=\xi(\xi^{2}+q^{2})$ refers to the phase function defined earlier. Lastly, we define the restriction norm spaces $X_{s,b}^{\delta}(\varphi)$ as the sets of all restrictions of $X_{s,b}(\varphi)$-functions to $[-\delta,\delta]\times\mathbb{R}\times\mathbb{T}$, where $\delta$ is a positive real number. These spaces will be equipped with the norm
@@
Since the phase function $\varphi$ wont vary throughout this paper, we will just write $X_{s,b}$ and $X_{s,b}^{\delta}$, respectively.
@@ 3. Linear and bilinear Strichartz estimates
Let $\varepsilon>0$ and $b>0$ be arbitrary. Furthermore, we define the bilinear Fourier multiplier $MP(\cdot,\cdot)$ by its Fourier transform
@@
We fix $\varepsilon>0$ and $b\coloneqq\frac{1}{2}+\varepsilon^{\prime}>0$. By employing Parseval’s identity, the Cauchy-Schwarz inequality and Fubini’s theorem, we obtain
@@
$c_{0}\lesssim\sum_{q_{1}\in\mathbb{Z}}\langle q_{1}\rangle^{-1-2\varepsilon}\int_{\mathbb{R}}\langle\tilde{\xi_{1}}\rangle^{-1-2\varepsilon^{\prime}}\mathrm{d}\tilde{\xi_{1}}<\infty.$
@@
    $\lVert MP(u,v)\rVert_{L_{txy}^{2}}\lesssim\lVert u\rVert_{X_{\frac{1}{2}+\varepsilon,\frac{1}{2}-\tilde{\varepsilon}}}\lVert v\rVert_{X_{0+\varepsilon,\frac{1}{2}-\tilde{\varepsilon}}},$ (5)
@@
and for fixed spatial frequencies $\xi$ or $q$, one can then invoke well-known estimates for the Schrödinger and Airy equations. This argument has been employed multiple times in the context of the Schrödinger propagator (see, e.g. [15, 14, 10]), whereas it appears to be new in the case of the Airy propagator. We make the described approach precise in the propositions that follow.
@@
We begin with the proof of (8). Taking (6) into account, the Sobolev embedding theorem for homogenous spaces, followed by Parseval’s identity in the $x$-variable and an application of Minkowski’s integral inequality, yields
@@
$=\Bigg{(}\int_{\mathbb{R}}\Bigg{(}\int_{\mathbb{T}}\int_{-T\lvert\xi\rvert}^{T\lvert\xi\rvert}\Big{\lvert}\Big{(}e^{i\tilde{t}\partial_{y}^{2}}\mathcal{F}_{x}u_{0}(\xi)\Big{)}(y)\Big{\rvert}^{4}\mathrm{d}\tilde{t}\mathrm{d}y\Bigg{)}^{\frac{1}{2}}\mathrm{d}\xi\Bigg{)}^{\frac{1}{2}},$
@@
Let $u_{0}$ and $v_{0}$ be functions such that $I_{y}^{\frac{1}{3}}u_{0}\in L_{xy}^{2}$ and $v_{0}\in L_{xy}^{2}$. Then the estimates
$\lVert I_{x}^{\frac{1}{6}}e^{-t\partial_{x}\Delta_{xy}}u_{0}\rVert_{L_{txy}^{6}}\lesssim\lVert I_{y}^{\frac{1}{3}}u_{0}\rVert_{L_{xy}^{2}}$ (14)
@@
hold. Furthermore, if $p\in[2,6]$, $b>\frac{1}{2}$, and $u$ is a function with $I_{y}^{\frac{1}{2}-\frac{1}{p}}u\in X_{0,(\frac{3}{2}-\frac{3}{p})b}$, then
$\lVert I_{x}^{\frac{1}{4}-\frac{1}{2p}}u\rVert_{L_{txy}^{p}}\lesssim_{b}\lVert I_{y}^{\frac{1}{2}-\frac{1}{p}}u\rVert_{X_{0,(\frac{3}{2}-\frac{3}{p})b}}$ (16)
@@
$\lVert I_{x}^{\frac{1}{6}}e^{-t\partial_{x}\Delta_{xy}}u_{0}\rVert_{L_{txy}^{6}}$ $\lesssim\lVert I_{y}^{\frac{1}{3}}I_{x}^{\frac{1}{6}}e^{-t\partial_{x}\Delta_{xy}}u_{0}\rVert_{L_{tx}^{6}L_{y}^{2}}$
$\sim\lVert\lvert q\rvert^{\frac{1}{3}}\Big{(}I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}u_{0}\Big{)}(x+q^{2}t)\rVert_{L_{tx}^{6}L_{q}^{2}}$
$\leq\lVert\lvert q\rvert^{\frac{1}{3}}\Big{(}I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}u_{0}\Big{)}(x+q^{2}t)\rVert_{L_{q}^{2}L_{tx}^{6}}$
$=\Bigg{(}\sum_{q\in\mathbb{Z}}\lvert q\rvert^{\frac{2}{3}}\Bigg{(}\int_{\mathbb{R}}\int_{\mathbb{R}}\Big{\lvert}\Big{(}I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}u_{0}(q)\Big{)}(x+q^{2}t)\Big{\rvert}^{6}\mathrm{d}x\mathrm{d}t\Bigg{)}^{\frac{1}{3}}\Bigg{)}^{\frac{1}{2}}.$
@@
... $=\Bigg{(}\sum_{q\in\mathbb{Z}}\lvert q\rvert^{\frac{2}{3}}\Bigg{(}\int_{\mathbb{R}}\int_{\mathbb{R}}\Big{\lvert}\Big{(}I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}u_{0}(q)\Big{)}(x)\Big{\rvert}^{6}\mathrm{d}x\mathrm{d}t\Bigg{)}^{\frac{1}{3}}\Bigg{)}^{\frac{1}{2}}$
$=\Bigg{(}\sum_{q\in\mathbb{Z}}\lvert q\rvert^{\frac{2}{3}}\lVert I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}(q)\rVert_{L_{tx}^{6}(\mathbb{R}\times\mathbb{R})}^{2}\Bigg{)}^{\frac{1}{2}},$
@@
... $\lesssim\Bigg{(}\sum_{q\in\mathbb{Z}}\lvert q\rvert^{\frac{2}{3}}\lVert\mathcal{F}_{y}u_{0}(q)\rVert_{L_{x}^{2}(\mathbb{R})}^{2}\Bigg{)}^{\frac{1}{2}}\sim\lVert I_{y}^{\frac{1}{3}}u_{0}\rVert_{L_{xy}^{2}},$
@@
$\lVert I_{x}^{\frac{1}{6}}u\rVert_{L_{txy}^{6}}\lesssim_{b}\lVert I_{y}^{\frac{1}{3}}u\rVert_{X_{0,b}},$ (17)
@@
$P_{\big{\{}\lvert\xi\rvert\leq\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}u\coloneqq\mathcal{F}_{xy}^{-1}\chi_{\big{\{}\lvert\xi\rvert\leq\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}\mathcal{F}_{xy}u$
@@
$P_{\big{\{}\lvert\xi\rvert>\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}u\coloneqq\mathcal{F}_{xy}^{-1}\chi_{\big{\{}\lvert\xi\rvert>\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}\mathcal{F}_{xy}u.$
@@
$\lVert u\rVert_{L_{Txy}^{p}}\leq\lVert P_{\big{\{}\lvert\xi\rvert\leq\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}u\rVert_{L_{Txy}^{p}}+\lVert P_{\big{\{}\lvert\xi\rvert>\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}u\rVert_{L_{Txy}^{p}},$
@@
$\chi_{\big{\{}\lvert\xi\rvert\leq\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}\langle\xi\rangle^{\frac{1}{2}-\frac{1}{p}}\langle q\rangle^{(\frac{3}{2}-\frac{3}{p})\varepsilon}\lesssim\langle q\rangle^{\frac{1}{3}-\frac{2}{3p}+(\frac{1}{2}-\frac{1}{p})\varepsilon},$
@@
$\lVert P_{\big{\{}\lvert\xi\rvert\leq\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}u\rVert_{L_{Txy}^{p}}$ $\lesssim_{T,\varepsilon,b}\lVert J_{x}^{\frac{1}{2}-\frac{1}{p}}J_{y}^{(\frac{3}{2}-\frac{3}{p})\varepsilon}P_{\big{\{}\lvert\xi\rvert\leq\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}u\rVert_{X_{0,(\frac{3}{2}-\frac{3}{p})b}}$
@@
$\chi_{\big{\{}\lvert\xi\rvert>\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}\lvert\xi\rvert^{-(\frac{1}{4}-\frac{1}{2p})}\lvert q\rvert^{\frac{1}{2}-\frac{1}{p}}\lesssim\langle q\rangle^{\frac{1}{3}-\frac{2}{3p}+(\frac{1}{2}-\frac{1}{p})\varepsilon},$
@@
$\lVert P_{\big{\{}\lvert\xi\rvert>\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}u\rVert_{L_{Txy}^{p}}$ $\lesssim_{T,\varepsilon,b}\lVert I_{x}^{-(\frac{1}{4}-\frac{1}{2p})}I_{y}^{\frac{1}{2}-\frac{1}{p}}P_{\big{\{}\lvert\xi\rvert>\lvert q\rvert^{\frac{2}{3}-2\varepsilon}\big{\}}}u\rVert_{X_{0,(\frac{3}{2}-\frac{3}{p})b}}$
@@
$\lvert B_{\tau,\xi,q}^{\mathrm{lin}}\rvert\lesssim\sum_{\lvert q_{1}+h\rvert\leq\big{(}\frac{3\xi(K-\lvert c\rvert)}{3\xi^{2}-q^{2}}\big{)}^{\frac{1}{2}}}b(q_{1})$

### v3

# Strichartz estimates for the generalized Zakharov-Kuznetsov equation on $\mathbb{R}\times\mathbb{T}$ and applications

@@ Abstract.
In this article, we address the Cauchy problem associated with the $k$-generalized Zakharov-Kuznetsov equation posed on $\mathbb{R}\times\mathbb{T}$. By establishing an almost optimal linear $L^{4}$-estimate, along with a family of bilinear refinements, we significantly lower the well-posedness threshold for all $k\geq 2$. In particular, we show that the modified Zakharov-Kuznetsov equation is locally well-posed in $H^{s}(\mathbb{R}\times\mathbb{T})$ for all $s>\frac{11}{24}$.
@@ 1. Introduction
$E_{\pm}(u(t))\coloneqq\frac{1}{2}\int_{\mathbb{R}\times\mathbb{T}}\lvert\nabla u(t,x,y)\rvert_{2}^{2}\pm\frac{2}{k+2}(u(t,x,y))^{k+2}\mathrm{d}(x,y)$
@@
In the case $k=1$, the foundational work was carried out by Linares, Pastor, and Saut [26], who established local well-posedness in $H^{s}(\mathbb{R}\times\mathbb{T})$ for all $s>\frac{3}{2}$. By developing a bilinear estimate that allows for a gain of up to $\frac{1}{2}$- derivatives, Molinet and Pilod [28] were subsequently able to prove global well-posedness for all $s\geq 1$. Based on a refinement of said estimate, Osawa [29] succeeded in lowering the threshold for local well-posedness to $s>\frac{9}{10}$, and, in joint work with Takaoka [30], further established global well-posedness for all $s>\frac{29}{31}$ by employing the $I$-method. More recently, Cao-Labora [5] developed a further refinement of said bilinear estimate, carefully adapted to the resonance function. By invoking geometric arguments, he was then able to improve the local well-posedness result to $s>\frac{3}{4}$ (or $s>\frac{1}{2}$ under an additional low-frequency condition), and in both cases established optimality up to the endpoint. For the case $k\geq 2$, Farah and Molinet [9] proved the linear $L^{4}$-estimate
@@
The development of increasingly sharp linear Strichartz estimates has become a central focus of current research concerning the well-posedness study of different problems posed on $\mathbb{R}\times\mathbb{T}$. Examples of such recent work include the contributions of Herr, Schippa, and Tzvetkov [18] on the semiperiodic dispersion-generalized KP-II equation; Corcho and Mallqui [6] on the semiperiodic NLS equation with fractional derivatives in the periodic component; and Başakoğlu, Sun, Tzvetkov, and Wang [1] on the semiperiodic hyperbolic NLS equations. In this spirit, the proofs of several new linear and bilinear estimates of Strichartz type will also serve as the starting point in this article. In particular, we wish to highlight the following linear $L^{4}$-estimate, which we will furthermore show to be optimal up to the loss of an $\varepsilon$-derivative.
@@
and let $s>s_{0}(k)$ be arbitrary. Then for every $u_{0}\in H^{s}(\mathbb{R}\times\mathbb{T})$, there exists a $\delta=\delta(\lVert u_{0}\rVert_{H^{s_{0}(k)+}(\mathbb{R}\times\mathbb{T})})>0$ and a unique solution
$u\in X_{s,\frac{1}{2}+}^{\delta}(\varphi)$
to (1). Moreover, for every $\tilde{\delta}\in(0,\delta)$, there exists a neighborhood $\mathcal{U}\subseteq H^{s}(\mathbb{R}\times\mathbb{T})$ of $u_{0}$ such that the flow map
@@
Subsequently, in the case $k=2$ (which corresponds to the modified Zakharov-Kuznetsov equation), we will carry out a detailed case-by-case analysis using the bilinear refinements of (2) and the resonance function, and push the local well-posedness threshold in this case down to $s>\frac{11}{24}$.
@@
Let $s>\frac{11}{24}$ be arbitrary. Then for every $u_{0}\in H^{s}(\mathbb{R}\times\mathbb{T})$, there exists a $\delta=\delta(\lVert u_{0}\rVert_{H^{\frac{11}{24}+}(\mathbb{R}\times\mathbb{T})})>0$ and a unique solution
@@ 2.1. Notation
If $A$ is a measurable subset of $\mathbb{R}\times\mathbb{Z}$, $\mathbb{R}\times\mathbb{R}\times\mathbb{Z}$ or $\mathbb{R}^{n}$, then $\lvert A\rvert$ will denote its Haar measure. For an arbitrary number $\in\mathbb{R}, x+x$+ will denote a number slightly greater than $x$, while xx- [^1] In the context of interpolation, the notation $\infty$- will be used to indicate a number that is very large, but still finite. will denote a number slightly smaller than $x$. If we have two real numbers, $x$ and $y$, then their maximum will be denoted by $x\lor y$ and their minimum by $x\land y$. If $x$ and $y$ are both positive, then $x\lesssim y$ will mean that there exists some constant $c>0$, such that $x\leq cy$ holds. In the case that the constant $c$ can be chosen particularly close to 0, we will write $x\ll y$ and if both $x\lesssim y$ and $y\lesssim x$ hold, we will write $x\sim y$. Furthermore, we will denote the absolute value of a real number $x$ by $\lvert x\rvert$, while for a vector $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, the Euclidean norm will be denoted by $\lvert x\rvert_{2}$. Using the Euclidean norm, we define the Japanese brackets as $\langle x\rangle\coloneqq(1+\lvert x\rvert_{2}^{2})^{\frac{1}{2}}$. For an admissible function $f=f(t,x,y)$, where $(t,x,y)\in\mathbb{R}^{2}\times\mathbb{T}$, its space-time Fourier transform will be denoted by $\hat{f}$, and its inverse space-time Fourier transform by $\check{f}$. If we are interested in the partial Fourier transform of such a function, we indicate this by using subscripts, e.g., in $\mathcal{F}_{x}f$ and $\mathcal{F}_{xy}^{-1}f$. We use the Fourier transform to define the Bessel and Riesz potential operators of order s-s, where $s$ is a real number:
@@
$R(\xi,q,\xi_{1},q_{1},\xi_{2},q_{2})\coloneqq\varphi(\xi,q)-\varphi(\xi_{1},q_{1})-\varphi(\xi_{2},q_{2})-\varphi(\xi-\xi_{1}-\xi_{2},q-q_{1}-q_{2}).$
@@ 2.2. Function spaces
where $\varphi(\xi,q)=\xi(\xi^{2}+q^{2})$ refers to the phase function defined earlier. Lastly, we define the restriction norm spaces $X_{s,b}^{\delta}(\varphi)$ as the sets of all restrictions of $X_{s,b}(\varphi)$-functions to $\delta,\delta]\times\mathbb{R}\times\mathbb{T}$, where $\delta$ is a positive real number. These spaces will be equipped with the norm
@@
Since the phase function $\varphi$ won’t vary throughout this paper, we will just write $X_{s,b}$ and $X_{s,b}^{\delta}$, respectively.
@@ 3. Linear and bilinear Strichartz estimates
Let $\varepsilon>0$ and $b>\frac{1}{2}$ be arbitrary. Furthermore, we define the bilinear Fourier multiplier $MP(\cdot,\cdot)$ by its Fourier transform
@@
We fix $\varepsilon>0$ and $b\coloneqq\frac{1}{2}+\varepsilon^{\prime}>\frac{1}{2}$. By employing Parseval’s identity, the Cauchy-Schwarz inequality and Fubini’s theorem, we obtain
@@
$c_{0}\lesssim\sum_{q_{1}\in\mathbb{Z}}\langle q_{1}\rangle^{-1-2\varepsilon}\int_{\mathbb{R}}\langle\tilde{\xi_{1}}\rangle^{-1-2\varepsilon^{\prime}}\mathrm{d}\tilde{\xi_{1}}$$\infty$.
@@
    $\lVert MP(u,v)\rVert_{L_{txy}^{2}}\lesssim\lVert u\rVert_{X_{\frac{1}{2}+\varepsilon,\frac{1}{2}-\tilde{\varepsilon}}}\lVert v\rVert_{X_{\varepsilon,\frac{1}{2}-\tilde{\varepsilon}}}$, (5)
@@
and for fixed spatial frequencies $\xi$ or $q$, one can then invoke well-known estimates for the Schrödinger and Airy equations. This argument has been employed multiple times in the context of the Schrödinger propagator (see, e.g. [17, 16, 12]), whereas it seems to be new in the case of the Airy propagator. We make the described approach precise in the propositions that follow.
@@
We begin with the proof of (8). Taking (6) into account, the Sobolev embedding theorem for homogeneous spaces, followed by Parseval’s identity in the $x$-variable and an application of Minkowski’s integral inequality, yields
@@
$\Bigg(\int_{\mathbb{R}}\Bigg(\int_{\mathbb{T}}\int_{-T\lvert\xi\rvert}^{T\lvert\xi\rvert}\Big\lvert\Big(e^{i\tilde{t}\partial_{y}^{2}}\mathcal{F}_{x}u_{0}(\xi)\Big)(y)\Big\rvert^{4}\mathrm{d}\tilde{t}\mathrm{d}y\Bigg)^{\frac{1}{2}}\mathrm{d}\xi\Bigg)^{\frac{1}{2}},$
@@
Let $u_{0}$ and $v_{0}$ be functions such that $J_{y}^{\frac{1}{3}}u_{0}\in L_{xy}^{2}$ and $v_{0}\in L_{xy}^{2}$. Then the estimates
$\lVert I_{x}^{\frac{1}{6}}e^{-t\partial_{x}\Delta_{xy}}u_{0}\rVert_{L_{txy}^{6}}\lesssim\lVert J_{y}^{\frac{1}{3}}u_{0}\rVert_{L_{xy}^{2}}$ (14)
@@
hold. [^2] The ”Airy endpoint estimate” will not be used in this paper, but it might be of independent interest for future applications. Furthermore, if $p\in[2,6]$, $b>\frac{1}{2}$, and $u$ is a function with $J_{y}^{\frac{1}{2}-\frac{1}{p}}u\in X_{0,(\frac{3}{2}-\frac{3}{p})b}$, then
$\lVert I_{x}^{\frac{1}{4}-\frac{1}{2p}}u\rVert_{L_{txy}^{p}}\lesssim_{b}\lVert J_{y}^{\frac{1}{2}-\frac{1}{p}}u\rVert_{X_{0,(\frac{3}{2}-\frac{3}{p})b}}$ (16)
@@
$\lVert I_{x}^{\frac{1}{6}}e^{-t\partial_{x}\Delta_{xy}}u_{0}\rVert_{L_{txy}^{6}}$ $\lesssim\lVert J_{y}^{\frac{1}{3}}I_{x}^{\frac{1}{6}}e^{-t\partial_{x}\Delta_{xy}}u_{0}\rVert_{L_{tx}^{6}L_{y}^{2}}$
$\sim\lVert\langle q\rangle^{\frac{1}{3}}\Big(I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}u_{0}\Big)(x+q^{2}t)\rVert_{L_{tx}^{6}L_{q}^{2}}$
$\leq\lVert\langle q\rangle^{\frac{1}{3}}\Big(I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}u_{0}\Big)(x+q^{2}t)\rVert_{L_{q}^{2}L_{tx}^{6}}$
$=\Bigg(\sum_{q\in\mathbb{Z}}\langle q\rangle^{\frac{2}{3}}\Bigg(\int_{\mathbb{R}}\int_{\mathbb{R}}\Big\lvert\Big(I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}u_{0}(q)\Big)(x+q^{2}t)\Big\rvert^{6}\mathrm{d}x\mathrm{d}t\Bigg)^{\frac{1}{3}}\Bigg)^{\frac{1}{2}}.$
@@
... $=\Bigg(\sum_{q\in\mathbb{Z}}\langle q\rangle^{\frac{2}{3}}\Bigg(\int_{\mathbb{R}}\int_{\mathbb{R}}\Big\lvert\Big(I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}u_{0}(q)\Big)(x)\Big\rvert^{6}\mathrm{d}x\mathrm{d}t\Bigg)^{\frac{1}{3}}\Bigg)^{\frac{1}{2}}$
$=\Bigg(\sum_{q\in\mathbb{Z}}\langle q\rangle^{\frac{2}{3}}\lVert I_{x}^{\frac{1}{6}}e^{-t\partial_{x}^{3}}\mathcal{F}_{y}(q)\rVert_{L_{tx}^{6}(\mathbb{R}\times\mathbb{R})}^{2}\Bigg)^{\frac{1}{2}},$
@@
... $\lesssim\Bigg(\sum_{q\in\mathbb{Z}}\langle q\rangle^{\frac{2}{3}}\lVert\mathcal{F}_{y}u_{0}(q)\rVert_{L_{x}^{2}(\mathbb{R})}^{2}\Bigg)^{\frac{1}{2}}\sim\lVert J_{y}^{\frac{1}{3}}u_{0}\rVert_{L_{xy}^{2}},$
@@
$\lVert I_{x}^{\frac{1}{6}}u\rVert_{L_{txy}^{6}}\lesssim_{b}\lVert J_{y}^{\frac{1}{3}}u\rVert_{X_{0,b}},$ (17)
@@
$P_{\big\{\lvert\xi\rvert\leq\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}u\coloneqq\mathcal{F}_{xy}^{-1}\chi_{\big\{\lvert\xi\rvert\leq\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}\mathcal{F}_{xy}u$
@@
$P_{\big\{\lvert\xi\rvert>\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}u\coloneqq\mathcal{F}_{xy}^{-1}\chi_{\big\{\lvert\xi\rvert>\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}\mathcal{F}_{xy}u.$
@@
$\lVert u\rVert_{L_{Txy}^{p}}\leq\lVert P_{\big\{\lvert\xi\rvert\leq\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}u\rVert_{L_{Txy}^{p}}+\lVert P_{\big\{\lvert\xi\rvert>\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}u\rVert_{L_{Txy}^{p}},$
@@
$\chi_{\big\{\lvert\xi\rvert\leq\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}\langle\xi\rangle^{\frac{1}{2}-\frac{1}{p}}\langle q\rangle^{(\frac{3}{2}-\frac{3}{p})\varepsilon}\lesssim\langle q\rangle^{\frac{1}{3}-\frac{2}{3p}+(\frac{1}{2}-\frac{1}{p})\varepsilon},$
@@
$\lVert P_{\big\{\lvert\xi\rvert\leq\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}u\rVert_{L_{Txy}^{p}}$ $\lesssim_{T,\varepsilon,b}\lVert J_{x}^{\frac{1}{2}-\frac{1}{p}}J_{y}^{(\frac{3}{2}-\frac{3}{p})\varepsilon}P_{\big\{\lvert\xi\rvert\leq\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}u\rVert_{X_{0,(\frac{3}{2}-\frac{3}{p})b}}$
@@
$\chi_{\big\{\lvert\xi\rvert>\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}\lvert\xi\rvert^{-(\frac{1}{4}-\frac{1}{2p})}\langle q\rangle^{\frac{1}{2}-\frac{1}{p}}\lesssim\langle q\rangle^{\frac{1}{3}-\frac{2}{3p}+(\frac{1}{2}-\frac{1}{p})\varepsilon},$
@@
$\lVert P_{\big\{\lvert\xi\rvert>\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}u\rVert_{L_{Txy}^{p}}$ $\lesssim_{T,\varepsilon,b}\lVert I_{x}^{-(\frac{1}{4}-\frac{1}{2p})}J_{y}^{\frac{1}{2}-\frac{1}{p}}P_{\big\{\lvert\xi\rvert>\langle q\rangle^{\frac{2}{3}-2\varepsilon}\big\}}u\rVert_{X_{0,(\frac{3}{2}-\frac{3}{p})b}}$
@@
$\lvert B_{\tau,\xi,q}^{\mathrm{lin}}\rvert\lesssim\sum_{\lvert q_{1}+h\rvert\leq\big(\frac{3\xi(K-\lvert c\rvert)}{3\xi^{2}-q^{2}}\big)^{\frac{1}{2}}}b(q_{1}).$